I need help with review problems from a worksheet. We aren't working in the text book yet. We just started a new semester. We are to simplify each product...The square root of 7/9 times the square root of 27/49. Also another problem of the square root of 3/16 times the square root of 4/25 times the square root of 5/36. Its been a while since Algebra I, and now I am in Geometry. He is reviewing first. Please help with these problems. Thank-you
Answer by stanbon(75887)

The square root of 7/9 times the square root of 27/49.
[sqrt(7/9)]=sqrt[27*7] / sqrt[49*9]
=[3sqrt 21]/[7*3
=[sqrt21]/7
-------------------


Also another problem of the square root of 3/16 times the square root of 4/25 times the square root of 5/36.
[sqrt(3/16)]
=(1/4)(sqrt3)*(2/5)*(1/6)sqrt5
=(1/60)sqrt15
------------
Cheers,
Stan H.